Evaluate.

negative square root of 16

cubed root of negative 27/64

Answer by checkley71(8403)

-SQRT16=-4 ANSWER.
------------------------
YOU ARE CORRECT. I MISSED THE (-) SIGN. SORRY.
CUBERT(-27/64)=CUBERT-(3*3*3)/(4*4*4)=-3/4 ANSWER.